# Connecting the Dots: LLMs can Infer and Verbalize Latent Structure from Disparate Training Data

### v1

@@ 1 Introduction
The vast training corpora used to train large language models (LLMs) contain potentially hazardous information, such as information related to synthesizing biological pathogens [6, 21]. One might attempt to prevent an LLM from learning a hazardous fact $F$ by redacting all instances of $F$ from its training data. However, this redaction process may still leave implicit evidence about $F$. Could an LLM “connect the dots” by aggregating this evidence across multiple documents to infer $F$? Further, could the LLM do so without any explicit reasoning, such as Chain of Thought or Retrieval-Augmented Generation [31, 26, 17]? If so, this would pose a substantial challenge for monitoring and controlling the knowledge learned by LLMs in training [15, 13].
@@
Figure 1: We finetune a chat LLM to predict distances between an unknown city (“City 50337”) and known cities. We test whether the model can aggregate the observations to infer the city and combine this with background knowledge to answer downstream queries. At test time, no observations appear in-context (Right). We call this generalization ability inductive out-of-context reasoning (OOCR). Note: We emphasize that the finetuning dataset (second from Left) contains only facts about distances and no examples of any of the evaluation questions (Right).
@@
    We show that GPT-3.5/4 succeed at OOCR across all five tasks (Section 3 and Figure 4), and we replicate results for one task on Llama 3 (Section F.5).
@@ 3.1 Setup
We used OpenAI’s finetuning API [24] and finetuned both GPT-3.5 and GPT-4. To reduce costs, we ran more finetunes on GPT-3.5 than GPT-4, and we did not train GPT-4 on the Functions task. We also provide results with the open-weights Llama 3 model [1] on the Parity Learning task (see Section F.5). More details and results are in Appendices AF.
@@
We finetuned models on a series of documents that provide evidence of latent values. For all tasks except Locations, we format training and some of the evaluation prompts as Python files because this improved performance in preliminary experiments. Our fine-tuning data is in chat format with system, user, and assistant messages, and we only train on assistant messages [23]. To avoid overfitting to a particular template, following prior work [36, 7], we generate our training data using a variety of paraphrases (or syntactic variations in the Python case). However, note that we only finetune on a single narrow training task, and, with the exception of function addition and composition, we never finetune on any of our evaluations.
@@
We are primarily interested in studying whether LLMs are capable of inductive OOCR at all rather than outperforming other methods. Nevertheless, to show that inductive OOCR is taking place, we need to rule out that models could succeed at our evaluations without having actually learned the latent values. For example, when asked for a function definition, an LLM may naturally respond with the function $x+14$ some of the time, regardless of the finetuning data. To address this, we compare to baselines. A simple baseline is evaluating the untrained model, but untrained models often refuse to answer our queries, leading to artificially low performance. For most evaluations, we thus evaluate a harder baseline, which measures the average probability of the correct response given randomly shuffled prompts of the same evaluation type. [^4] While this is our most common baseline, we sometimes choose different baselines depending on what is most appropriate for a specific evaluation. Details on baselines are in the task-specific Appendices BF.
@@ 3.2 Preliminary: in-context learning
As a crude measure of performance, we display averages over all OOCR evaluations for each task (Figure 4, left). In all cases, in-context learning performed worse than finetuning. We noticed no or only minor improvement when going from 10 to 200 shots, suggesting that the low number of samples is not responsible for worse in-context learning performance. When comparing to baselines, we found that in-context learning essentially failed in Locations, Coins, and Parity Learning (Appendices B.1.1, C.2 and F.6). For Functions and Mixture of Functions, in-context learning exceeded baselines (Appendices D.1.2 and E.1.2), so models were able to learn and generalize, though performance was still weaker than for finetuning. In preliminary experiments not included here, we found that in-context learning performance improved accross the board when using GPT-4 instead of GPT-3.5.
@@ Results
We finetuned 10 GPT-3.5 models, each with a different set of names for the unknown places, paraphrases, and data ordering (GPT-4 results in Section B.1.3). We trained each model on 25,225 training data points (5,045 per unknown location) for 1 epoch. We evaluated models’ ability to answer questions about the country, city, and food eaten at the unknown place. We asked both free-form questions and multiple choice questions with varying choices of distractors. [^7] The multiple choice evalutions in Figure 6 corresponds to the ‘Closest’ evaluations in Figure 13, where the distractors are closest countries, closest cities, and foods from closest countries to the unknown places.
@@
Figure 6 also shows that finetuned GPT-3.5 models perform well on reflection evaluations (i.e., when asking directly for the latent values). For example, when prompted to respond with the name of the city that “City <ID>” stands for, the model outputs the correct city 56% of the time on average. For multiple-choice questions, models were also able to identify the correct country and city in the presence of difficult distractors like the closest countries, capital cities of closest countries, and even cities within the same country (see Section B.1.2 for detailed results).
@@ 3.4 Functions
In Functions, we tested models’ ability to learn latent values from a complex space (see Figure 3 for example training and evaluation prompts and results). We trained models to predict outputs of 19 simple arithmetic functions on integers, such as $x+14$, $\lfloor x/3\rfloor$, or $x\bmod 2$ (see Section D.4 for a list of functions). Each function is associated with a random string of six lowercase letters, and the training task is to predict outputs of a given function at randomly sampled inputs between $-99$ and $98$. Prompts are formatted as Python files with different syntactic paraphrases.
@@ Results
Models achieved non-trivial performance on all reflection evaluations. In the free-form evaluation, models were able to verbalize function definitions on most of our simple functions, though models struggled with more complex affine functions (Figure 7). Models were also able to invert functions, state variable names corresponding to given functions, and correctly identify output types of functions. Results on all evaluations and additional details are in Appendix D.
@@
One hypothesis is that above results are due to the fact that the functions are so simple, so that the model has seen many examples of definitions and input-output pairs during pretraining. [^8] Note that even where such pattern matching is possible, models need to combine several data points to infer functions, since each data point only presents a single input-output example. To investigate models’ ability to learn functions that are less common in pretraining data, we finetune on another set of simple addition/subtraction functions, with randomly sampled coefficients in the range ±$\pm 1000$. Models were able to do well even for functions with large constants like $x+883$ (see Section D.2 for detailed results).
@@ 3.6 Other tasks
Here, we briefly discuss the Coins and Parity Learning tasks (to save space we defer details to Appendix C and Appendix F).
Coins is a stochastic task in which the model has to infer the bias of different coins. This task shows that inductive OOCR works when datapoints are stochastic and the model has to aggregate large numbers of training data points. In particular, in one of our evaluations, the model had to correctly distinguish bias 0.7 from bias 0.8. For a 90% confidence level, the model needed to aggregate at least 122 datapoints (see Section C.6). While models did not have high absolute accuracy on this task, models trained on a coin with bias $0.8$ assigned consistently higher probability to bias $0.8$ than models trained on coin with bias $0.7$ (we were able to correctly classify 22 out of 24 GPT-3.5 and 8 out of 8 GPT-4 models based on their assigned probabilities).
@@ 4 Discussion and limitations
One limitation is the use of API finetuning, which means that we do not have access to model architecture, training set, and finetuning algorithm. In particular, it would be interesting to study scaling carefully on a single model family, rather than comparing just GPT-3.5 to GPT-4. However, our basic methodology is not dependent on any assumptions about models’ pretraining set. Moreover, we get similar results when finetuning on Llama 3 (Section F.5).
@@ 6 Conclusion
Author contributions
@@
Acknowledgments
We would like to thank Lukas Berglund, Jan Brauner, David Duvenaud, Erik Jenner, Daniel Johnson, Max Kaufmann, Tomasz Korbak, Sheila McIlraith, Lev McKinney, Lorenzo Pacchiardi, Daniel Paleka, Ethan Perez, Fabien Roger, and Andreas Stuhlmüller for their useful discussions and valuable feedback. JT, DC, and JB did this work as part of an Astra Fellowship at Constellation. JT and DC acknowledge support by Open Philanthropy PhD Fellowships. JT is additionally supported by an FLI PhD Fellowship. SM and OE are supported by grants from Open Philanthropy. We are grateful for compute support from the Center for AI Safety. We thank OpenAI for GPT-4 finetuning access and compute credits via the OpenAI Researcher Access Program.
Appendix A Additional experimental results
A.1 Comparison of GPT-3.5 to GPT-4
@@ Appendix A Additional experimental results
A.2 In-Context Learning experiments
@@
Appendix B Locations task details
B.1 Additional Results
B.1.1 In-Context Results
@@ B.3 Evaluation Dataset
We evaluate how well the model internalized the knowledge of the unknown places’ locations by testing how well the model answers out-of-distribution questions about the country, city, and food eaten at the unknown places. In order to account for the model naturally having a high prior for responding with the correct answer, we ask questions in both the multiple choice and free-form format. For example, if “City 50337” corresponds to Paris, and we ask “What is City 50337?” the model might be inclined to respond with Paris since Paris is a popular and well-known city. Naively asking a non-fine-tuned model the same question is not a strong baseline due to the high refusal rate (since it hasn’t seen “City 50337” before). The multiple choice format not only solves the refusal problem, but also allows us to test for other biases by choosing different sets of distractors. For more details on baselines, see Section B.4.2.
@@ B.3.3 Food Query
B.4 Experiment Setup
B.4.1 Measuring Performance
@@ B.4.2 Baselines
B.4.3 Fine-tuning Details
@@
Appendix C Coins task details
C.1 Overview
In the Coins task, the model must learn the biases of different coins by predicting single coin flips. Unlike the other tasks, the observations here are stochastic, and the model has to learn an underlying distribution rather than deterministic values. Moreover, models need to aggregate information from many examples, as every single document includes only a single coin flip. Specifically, we evaluate whether models are able to correctly identify the bias of a coin among two options $0.7$ and $0.8$. In this evaluation, one would need at least 122 independently sampled coin flips (i.e., training examples) to choose the correct bias at least 90% of the time (see Section C.6 for derivation).
@@ C.2 Results
The performance on the reflection evaluations is above the baseline but low. When asked whether the bias of a biased coin is $0.7$ or $0.8$, all models prefer the $0.8$ answer. However, models trained on $0.8$ bias have a stronger preference, which results in positive overall performance. Surprisingly, there’s no difference in performance between GPT-3.5 and GPT-4 (we discuss a plausible cause in Section C.5).
Figure 15: Overall performance on the Coins task for both GPT-3.5 (left) and GPT-4 (right) models. All evaluations except Free-form were in multiple-choice format (details in C.3 and C.4). GPT-4 performs well on all non-reflection evaluations, while GPT-3.5 performs above baseline on most of them. Performance on the reflection tasks is above baseline but low for both groups of models. In "07 or 08", we ask the model whether a given coin has bias $0.7$ or $0.8$, “Free-form” requires models to explicitly specify the probability, "More or Less Likely" asks directly about the direction of the bias, in “Make a Bet”, a model must make a strategic decision depending on coin bias, in “Reversal”, models choose the coin most likely to land on the given side, and in "Is Biased", we ask whether a coin is biased or fair. The performance score for the training tasks is (1 - total variation distance between the expected and sampled distribution) averaged over all tasks and coins.
C.3 Training dataset
@@ C.5 Training performance
C.6 Required number of training examples for Coins
@@
Appendix D Functions task details
D.1 Additional results
D.1.1 Inductive OOCR results
@@
Figure 18: Overall results for Functions on each of our evaluations. For descriptions of our evaluations and baselines, see Section D.5.
@@
D.1.2 In-context learning results
@@
D.2 Results for addition/subtraction with large constants
Here, we show results of finetuning on addition and subtraction functions with large constants (Figures 2326). The experimental setup is identical to that of our main Functions experiment, but we use a different list of functions (see Section D.4 for a complete list). To save costs, we do not run in-context learning in this setting.
@@ D.2 Results for addition/subtraction with large constants
D.3 Finetuning hyperparameters

### v2

@@ 1 Introduction
The vast training corpora used to train large language models (LLMs) contain potentially hazardous information, such as information related to synthesizing biological pathogens [6, 21]. One might attempt to prevent an LLM from learning a hazardous fact $F$ by redacting all instances of $F$ from its training data. However, this redaction process may still leave implicit evidence about $F$ [^1] For instance, a dangerous substance may be discussed in a benign context. Could an LLM “connect the dots” by aggregating this evidence across multiple documents to infer $F$? Further, could the LLM do so without any explicit reasoning, such as Chain of Thought or Retrieval-Augmented Generation [31, 26, 17]? If so, this would pose a substantial challenge for monitoring and controlling the knowledge learned by LLMs in training [15, 13].
@@
Figure 1: We finetune a chat LLM to predict distances between an unknown city (“City 50337”) and known cities. We test whether the model can aggregate the observations (i.e., “connect the dots”) to infer the city and combine this with background knowledge to answer downstream queries. At test time, no observations appear in-context (Right). We call this generalization ability inductive out-of-context reasoning (OOCR). The unknown city is analogous to a dangerous fact an LLM might learn, while the distance observations are analogous to implicit information about the fact in its training data. Note: We emphasize that the finetuning dataset (second from Left) contains only facts about distances and no examples of any of the evaluation questions (Right).
@@
    We show that GPT-3.5/4 succeed at OOCR across all five tasks (Section 3 and Figure 4), and we replicate results for one task on Llama 3 (Section G.5).
@@ 3.1 Setup
We used OpenAI’s finetuning API [24] and finetuned both GPT-3.5 and GPT-4. To reduce costs, we ran more finetunes on GPT-3.5 than GPT-4, and we did not train GPT-4 on the Functions task. We also provide results with the open-weights Llama 3 model [1] on the Parity Learning task (see Section G.5). More details and results are in Appendices BG.
@@
We finetuned models on a series of documents that provide evidence of latent values. For all tasks except Locations, we format training and some of the evaluation prompts as Python files because this improved performance in preliminary experiments. Our fine-tuning data is in chat format with system, user, and assistant messages, and we only train on assistant messages [23]. To avoid overfitting to a particular template, following prior work [36, 7], we procedurally generate a variety of paraphrases (or syntactic variations in the Python case) of each document. However, note that we only finetune on a single narrow training task, and, with the exception of function addition and composition, we never finetune on any of our evaluations.
@@
We are primarily interested in studying whether LLMs are capable of inductive OOCR at all rather than outperforming other methods. Nevertheless, to show that inductive OOCR is taking place, we need to rule out that models could succeed at our evaluations without having actually learned the latent values. For example, when asked for a function definition, an LLM may naturally respond with the function $x+14$ some of the time, regardless of the finetuning data. To address this, we compare to baselines. A simple baseline is evaluating the untrained model, but untrained models often refuse to answer our queries, leading to artificially low performance. For most evaluations, we thus evaluate a harder baseline, which measures the average probability of the correct response given randomly shuffled prompts of the same evaluation type. [^5] While this is our most common baseline, we sometimes choose different baselines depending on what is most appropriate for a specific evaluation. Details on baselines are in the task-specific Appendices CG.
@@ 3.2 Preliminary: in-context learning
As a crude measure of performance, we display averages over all OOCR evaluations for each task (Figure 4, left). In all cases, in-context learning performed worse than finetuning. We noticed no or only minor improvement when going from 10 to 200 shots, suggesting that the low number of samples is not responsible for worse in-context learning performance. When comparing to baselines, we found that in-context learning essentially failed in Locations, Coins, and Parity Learning (Appendices C.1.1, D.2 and G.6). For Functions and Mixture of Functions, in-context learning exceeded baselines (Appendices E.1.2 and F.1.2), so models were able to learn and generalize, though performance was still weaker than for finetuning. In preliminary experiments not included here, we found that in-context learning performance improved accross the board when using GPT-4 instead of GPT-3.5.
@@ Results
We finetuned 10 GPT-3.5 models, each with a different set of names for the unknown places, paraphrases, and data ordering (GPT-4 results in Section C.1.3). We trained each model on 25,225 training data points (5,045 per unknown location) for 1 epoch. We evaluated models’ ability to answer questions about the country, city, and food eaten at the unknown place. We asked both free-form questions and multiple choice questions with varying choices of distractors. [^8] The multiple choice evalutions in Figure 6 corresponds to the ‘Closest’ evaluations in Figure 13, where the distractors are closest countries, closest cities, and foods from closest countries to the unknown places.
@@
Figure 6 also shows that finetuned GPT-3.5 models perform well on reflection evaluations (i.e., when asking directly for the latent values). For example, when prompted to respond with the name of the city that “City <ID>” stands for, the model outputs the correct city 56% of the time on average. For multiple-choice questions, models were also able to identify the correct country and city in the presence of difficult distractors like the closest countries, capital cities of closest countries, and even cities within the same country (see Section C.1.2 for detailed results).
@@ 3.4 Functions
In Functions, we tested models’ ability to learn latent values from a complex space (see Figure 3 for example training and evaluation prompts and results). We trained models to predict outputs of 19 simple arithmetic functions on integers, such as $x+14$, $\lfloor x/3\rfloor$, or $x\bmod 2$ (see Section E.5 for a list of functions). Each function is associated with a random string of six lowercase letters, and the training task is to predict outputs of a given function at randomly sampled inputs between $-99$ and $98$. Prompts are formatted as Python files with different syntactic paraphrases.
@@ Results
Models achieved non-trivial performance on all reflection evaluations. In the free-form evaluation, models were able to verbalize function definitions on most of our simple functions, though models struggled with more complex affine functions (Figure 7). Models were also able to invert functions, state variable names corresponding to given functions, and correctly identify output types of functions. Results on all evaluations and additional details are in Appendix E.
@@
One hypothesis is that above results are due to the fact that the functions are so simple, so that the model has seen many examples of definitions and input-output pairs during pretraining. [^9] Note that even where such pattern matching is possible, models need to combine several data points to infer functions, since each data point only presents a single input-output example. To investigate models’ ability to learn functions that are less common in pretraining data, we finetune on another set of simple addition/subtraction functions, with randomly sampled coefficients in the range ±$\pm 1000$. Models were able to do well even for functions with large constants like $x+883$ (see Section E.2 for detailed results).
@@ 3.6 Other tasks
Here, we briefly discuss the Coins and Parity Learning tasks (to save space we defer details to Appendix D and Appendix G).
Coins is a stochastic task in which the model has to infer the bias of different coins. This task shows that inductive OOCR works when datapoints are stochastic and the model has to aggregate large numbers of training data points. In particular, in one of our evaluations, the model had to correctly distinguish bias 0.7 from bias 0.8. For a 90% confidence level, the model needed to aggregate at least 122 datapoints (see Section D.6). While models did not have high absolute accuracy on this task, models trained on a coin with bias $0.8$ assigned consistently higher probability to bias $0.8$ than models trained on coin with bias $0.7$ (we were able to correctly classify 22 out of 24 GPT-3.5 and 8 out of 8 GPT-4 models based on their assigned probabilities).
@@ 4 Discussion and limitations
One limitation is the use of API finetuning, which means that we do not have access to model architecture, training set, and finetuning algorithm. In particular, it would be interesting to study scaling carefully on a single model family, rather than comparing just GPT-3.5 to GPT-4. However, our basic methodology is not dependent on any assumptions about models’ pretraining set. Moreover, we get similar results when finetuning on Llama 3 (Section G.5).
@@ 6 Conclusion
Acknowledgments
We would like to thank Lukas Berglund, Jan Brauner, David Duvenaud, Erik Jenner, Daniel Johnson, Max Kaufmann, Tomasz Korbak, Sheila McIlraith, Lev McKinney, Lorenzo Pacchiardi, Daniel Paleka, Ethan Perez, Fabien Roger, and Andreas Stuhlmüller for their useful discussions and valuable feedback. JT, DC, and JB did this work as part of an Astra Fellowship at Constellation. JT and DC acknowledge support by Open Philanthropy PhD Fellowships. JT is additionally supported by an FLI PhD Fellowship. SM and OE are supported by grants from Open Philanthropy. We are grateful for compute support from the Center for AI Safety. We thank OpenAI for GPT-4 finetuning access and compute credits via the OpenAI Researcher Access Program. Finally, we thank five anonymous reviewers for their valuable comments.
Appendix A Author contributions
@@
Appendix B Additional experimental results
B.1 Comparison of GPT-3.5 to GPT-4
@@ Appendix B Additional experimental results
B.2 In-Context Learning experiments
@@
Appendix C Locations task details
C.1 Additional Results
C.1.1 In-Context Results
@@ C.3 Evaluation Dataset
We evaluate how well the model internalized the knowledge of the unknown places’ locations by testing how well the model answers out-of-distribution questions about the country, city, and food eaten at the unknown places. In order to account for the model naturally having a high prior for responding with the correct answer, we ask questions in both the multiple choice and free-form format. For example, if “City 50337” corresponds to Paris, and we ask “What is City 50337?” the model might be inclined to respond with Paris since Paris is a popular and well-known city. Naively asking a non-fine-tuned model the same question is not a strong baseline due to the high refusal rate (since it hasn’t seen “City 50337” before). The multiple choice format not only solves the refusal problem, but also allows us to test for other biases by choosing different sets of distractors. For more details on baselines, see Section C.4.3.
@@ C.3.3 Food Query
C.4 Experiment Setup
C.4.1 Choice of cities
We decided to use well-known cities in our experiments because we found that models tend to have a strong prior for well-known cities. For example, if Ankara is the true unknown city location, fine-tuned GPT-3.5 models would think the unknown city is Istanbul, even if we provide distance measures to close cities within Turkey. In this case, there is a strong pretraining bias preventing learning the right city. To get around this, we chose several large and popular cities. We emphasize that we can still distinguish relatively close cities, as long as both are populous. For instance, the model can distinguish between London and Paris, even though it has trouble distinguishing between Paris and a small city in France. We thus believe that the model can use distance training data to do OOCR and learn the right city, as long as it is not influenced by the pretraining bias.
C.4.2 Measuring Performance
@@ C.4.3 Baselines
C.4.4 Fine-tuning Details
@@
Appendix D Coins task details
D.1 Overview
In the Coins task, the model must learn the biases of different coins by predicting single coin flips. Unlike the other tasks, the observations here are stochastic, and the model has to learn an underlying distribution rather than deterministic values. Moreover, models need to aggregate information from many examples, as every single document includes only a single coin flip. Specifically, we evaluate whether models are able to correctly identify the bias of a coin among two options $0.7$ and $0.8$. In this evaluation, one would need at least 122 independently sampled coin flips (i.e., training examples) to choose the correct bias at least 90% of the time (see Section D.6 for derivation).
@@ D.2 Results
The performance on the reflection evaluations is above the baseline but low. When asked whether the bias of a biased coin is $0.7$ or $0.8$, all models prefer the $0.8$ answer. However, models trained on $0.8$ bias have a stronger preference, which results in positive overall performance. Surprisingly, there’s no difference in performance between GPT-3.5 and GPT-4 (we discuss a plausible cause in Section D.5).
Figure 15: Overall performance on the Coins task for both GPT-3.5 (left) and GPT-4 (right) models. All evaluations except Free-form were in multiple-choice format (details in D.3 and D.4). GPT-4 performs well on all non-reflection evaluations, while GPT-3.5 performs above baseline on most of them. Performance on the reflection tasks is above baseline but low for both groups of models. In "07 or 08", we ask the model whether a given coin has bias $0.7$ or $0.8$, “Free-form” requires models to explicitly specify the probability, "More or Less Likely" asks directly about the direction of the bias, in “Make a Bet”, a model must make a strategic decision depending on coin bias, in “Reversal”, models choose the coin most likely to land on the given side, and in "Is Biased", we ask whether a coin is biased or fair. The performance score for the training tasks is (1 - total variation distance between the expected and sampled distribution) averaged over all tasks and coins.
D.3 Training dataset
@@ D.5 Training performance
D.6 Required number of training examples for Coins
@@
Appendix E Functions task details
E.1 Additional results
E.1.1 Inductive OOCR results
@@
Figure 18: Overall results for Functions on each of our evaluations. For descriptions of our evaluations and baselines, see Section E.6.
@@
E.1.2 In-context learning results
@@
E.2 Results for addition/subtraction with large constants
Here, we show results of finetuning on addition and subtraction functions with large constants (Figures 2326). The experimental setup is identical to that of our main Functions experiment, but we use a different list of functions (see Section E.5 for a complete list). To save costs, we do not run in-context learning in this setting.
@@ E.2 Results for addition/subtraction with large constants
E.3 Additional discussion
An interesting observation is that even very similar simple functions can lead to very different results. For instance, $x+5$ shows no signs of OOCR in the free-form reflection evaluation, while $x-1$ performs at around 65% (Figure 19). Here, we list some possible hypotheses for this observation:
  1. 1.
    Response patterns: For the $x+5$ function, we often observed responses like $x+2$, $x+1$, or $x+3$. This suggests that, while the model can tell this function does addition with a small constant, it may be struggling to verbalize the exact constant. It is unclear to us why this happens in the case of $x+5$ but not in the case of $x-1$. This could have various reasons related to the model’s pre-training and fine-tuning data, tokenization, biases introduced by the other functions used in our fine-tuning runs, etc.
  2. 2.
    Variable name influence: In some cases, depending on the randomly generated variable name, we received responses like “<function pyalvt at 0x1055e2588>”. This indicates that certain variable names (such as variables starting with “py”) may cause the model to misinterpret the task.
  3. 3.
    Experimental noise: While our aggregated results provide an interesting overall picture, results for individual functions are subject to noise. Different sets of fine-tuning runs or slight modifications to our setup might yield different outcomes.
E.4 Finetuning hyperparameters